# PixFoundation: Are We Heading in the Right Direction with Pixel-level Vision Foundation Models?

### v2

@@ 3.2 A Pixel-level MLLMs study
| Image | Prompt | Concept Category | Noun Phrase | Output |
| 1 | Identify the butterfly’s wings in the scene. | Color & Appearance | orange wings | In the image, there is a butterfly with orange wings. |
| 3 | Identify the flame of the match in the scene. | Location & Position | the top | The flame of the match is located at the top of the image, surrounded by darkness. |
| 6 | Identify the dog’s face in the scene. | Color & Appearance | a black and white dog | The dog’s face in the scene is a black and white dog with a black nose. |
| 161 | Identify the minute hand of the clock in the scene. | Location & Position | the 12 o’clock position | The minute hand of the clock in the scene is located at the 12 o’clock position. |
[5pt]
1
[5pt]
3
[5pt]
6
[5pt]
161
@@ 4.1 Experimental Setup
Evaluation benchmarks, protocols and metrics. PixMMVP is composed of 300 images paired with questions, choices, referring expressions and segmentation masks, while PixCV-Bench has 1,438 images with their corresponding annotations similarly. On each benchmark we evaluate the VQA and visual grounding capabilities following three probing techniques and reporting their metrics. The first probing is to evaluate the VQA ability, where the accuracy is computed using GPT-4o following [19] as, . If the model generates a segmentation without explicitly asking it to, it is evaluated with respect to the ground-truth referring segmentation in terms of mean intersection over union as, . The second probing prompts the model to identify the referred expression then evaluates the mean intersection over union reported as, . The third probing following [18] instructs the model to generate a single option letter and evaluate the accuracy directly without GPT-4o, reported as, . There is a need for the first probing since some of the recent pixel-level MLLMs face challenges in following instructions. We evaluate the score of each model, , which is the harmonic mean across the maximum of both pixel-level visual grounding and VQA,
We mainly focus on evaluating four state-of-the-art pixel-level MLLMs; LISA [9], GLAMM [16], OMG-Llava [27] and Llava-G [26]. For GLAMM we use two variants; the original model (GLAMM) and the one fine-tuned for region-level captioning, (GLAMM-RegCap). For details on the models weights, refer to App. A.
Baselines and upper bound implementation details. We evaluate: (i) the attend and segment (a+s), (ii) the oracle selection relying on the highest intersection over union in selecting the correctly predicted masks (PixFoundation), and (iii) the automatic selection, (PixFoundation). These are implemented on top of three base MLLMs which are, Llava 1.5 (7B, 13B) [12] and Cambrian-1(8B) [18]. The automatic selection is implemented using GPT-4o, refer to App. A for more details.
@@ 4.2 Are the current pixel-level MLLMs heading in the right direction?
In order to answer this, we evaluate each of these pixel-level MLLMs VQA capabilities in challenging tasks. Additionally, we evaluate their ability to visually ground the objects of interest in these questions. Table 1 shows the results on the challenging PixMMVP and PixCV-Bench. From the accuracy of VQA, MLLMs that are not trained with pixel-level grounding surpass their pixel-level counterpart with up to 16%. The best in pixel-level MLLMs score in this aspect is GLAMM-RegCap [27] yet it has degraded ability to generate segmentation. On the other hand, when looking at pixel-level visual grounding we find the best model, GLAMM [16], has a weak ability in VQA or following instructions. Moreover, it shows LISA and Llava-G are mostly incapable of following the instruction to output the option letter reported in, . OMG-Llava strikes the right balance in both VQA and pixel-level grounding with the highest score, , within pixel-level MLLMs. However, looking at the bottom three rows, the oracle, PixFoundation, confirms that MLLMs that were never trained with pixel-level grounding have the correct grounding within their learned attention maps yielding higher score, , than OMG-Llava. Qualitative comparison is provided in Fig. 5 and additional examples are in App. B. Additionally, the automatic baseline outperforms OMG-Llava in PixMMVP with a considerable margin and is on-par to it on PixCV-Bench. Furthermore, the attend and segment baseline [2] lags behind our automatic method with more than 10% in both datasets. Refer to App. C for a failure case analysis.
Finally, we evaluate whether the failures of these MLLMs occur in visual grounding, VQA or both. Figure 6 shows the frequency of failures per category, where the majority stem from failures in both especially in the pixel-level MLLMs. While the vanilla MLLMs perform better in the VQA and tend to fail more in the grounding task.
Summary. In summary, pixel-level grounding supervision degrades MLLMs ability in VQA and sometimes even their generalization in grounding. We show that MLLMs trained with pixel-level supervision lag behind vanilla MLLMs using simple mechanisms to extract grounding, and the oracle indicates there is an opportunity to improve this. Moreover, we show that grounding might not coincide with the noun phrase most similar to the referred expression, where our oracle upper bound and automatic baseline are both surpassing the attend and segment.
@@ 4.3 When does grounding emerge in MLLMs?
When - location. Taking into account the powerful performance of the oracle upper bound, it begs the important question of when grounding emerges. We start by looking at when it emerges in terms of the location within the output text. We analyze the word/phrase location with respect to the full output text in terms of a percentage from its total length, (i.e., 0% means the beginning of the text). Accordingly, Fig. 7(a) shows the location percentages histogram, binned at 10%, for the three base MLLMs reporting the oracle selection and evaluating on PixMMVP benchmark using the second probing. In the Llava 1.5 variants, the highest grounding emerges at the last 40% of the output text, while for Cambrian it emerges at the last 60%.
When - concept. For the second analysis we look into the concept category that the correct output word/phrase corresponds to. The previous assumption in other works is that grounding emerges in the exact noun/noun phrase of the object of interest. Except our analysis confirms that this is not necessarily the case. We take the correct noun/noun phrase where the grounding emerges based on the oracle from all the three variants, then we pass it to GPT-4o to request a grouping of these concepts. It resulted into six main groups, which are: (i) Color and appearance, (ii) location and position, (iii) object parts, (iv) context and setting, (v) objects and entities, and (vi) State. We then prompt for each of the noun/noun phrase, GPT-4o, to categorize it within these six categories. The histogram of the occurrences of these concept categories is shown in Fig. 7(b). It clearly conveys that in certain scenarios the correct output when grounding emerges can be describing the position or the color of the object of interest not necessarily the exact ground-truth referring expression. Fig. 8 shows qualitative examples of these scenarios from PixMMVP using the second probing. Additional results are provided in App. B.
Random vs. best. All the results of our baselines and our findings hinge on the fact that we are using the maximum attention per output noun phrase to prompt SAM for the segmentation mask. Nonetheless, as a lower bound analysis, we evaluate the performance if we use a random point as prompt to SAM instead. For fair comparison, we generate random points with the count of output masks that the oracle has to select among (i.e., the number of the output noun phrases). We conduct this ablation on PixMMVP using Llava 1.5 (7B) base MLLM, with random point prompts followed by the oracle selection among their SAM masks. Figure 9 shows that this random + oracle baseline is a strong baseline, yet it lags behind the correct one using the maximum point (i.e., First) with around 12%. More importantly, we confirm the stability of the results if we select the second best or third best attention (i.e., Second and Third), which are on-par to the maximum point. Thus, even with the oracle selection, using the wrong point prompts lags with a considerable margin behind using the correct ones.
Summary. In summary, we found that highest frequent emergence of grounding is in the last 40-60% of the output text on PixMMVP which might indicate that the reasoning of the model on this challenging task to provide a response impacts when the correct visual grounding can occur. More importantly, we show that grounding in MLLMs can emerge in the noun phrase that corresponds to color, position or other characteristics of the object of interest and not necessarily the exact referring expression.
@@ 5 Conclusion
We proposed two benchmarks showing that pixel-level MLLMs degraded the ability in VQA and even grounding of fine-grained objects. Thus, our results questioned whether we are heading in the right direction with these models. Additionally, we provide powerful baselines with improved scores without training for pixel-level grounding.
Impact Statement
@@ Appendix A Additional implementation details
In this section, we cover additional details about our proposed datasets and the implementation of the evaluation setup and baselines. We also refer to the output from the questions of the three probing techniques in the supplementary material for all the studied models.
@@
Datasets. Our proposed datasets, PixMMVP and PixCV-Bench, are composed of ground-truth referring expressions describing the object of interest in the respective question and its segmentation mask. We show in Fig. 10 examples of these ground-truth annotations for both datasets. It shows the challenging scenarios in pixel-level visual grounding which is strongly tied to the visual question answering task, since an integral part of answering these questions requires the grounding of the object/s of interest.
Models. We also detail the model checkpoints we use for the four pixel-level MLLMs and their variants, retrieved from HuggingFace [22] in Table 2. These also include the model checkpoints used for the base MLLMs that were not trained with pixel-level visual grounding. It is worth noting, that for GLAMM we use two variants (FullScope and RegCap) since their base model (i.e., FullScope) has low performance in the visual question answering task. As such we use the other variant for GLAMM that was fine-tuned for region-level captioning using RefCOCOg dataset. Furthermore, we provide details on the oracle selection mechanism, we discard the cases where the ground-truth segmentation is all background in the when analysis, since there is no ground-truth grounding to evaluate against. These occur in few cases in PixMMVP.

### v4

@@ 3.3 Interpretability mechanism
For our interpretability mechanism, we build upon the previous pipeline and provide an oracle upper bound and an automatic method, which we call PixFoundation. We propose an additional mask selection by using MLLM as a judge on the different segmentation masks corresponding to the output tokens, to select the best mask and its respective noun phrase. Specifically, for each noun phrase in the output, we average the attention maps, , for each respective token, . Afterwards, we extract the maximum attention point, prompt SAM with it, highlight the generated mask in the image and ask the MLLM to select the highlighted image that best describes the queried referring expression from all the masks and their respective noun phrases.
Furthermore, we consider that the point prompt ambiguity is crucial in identifying fine-grained details; as such, we allow for our mask selection to consider three masks from SAM for each point and its respective noun phrase. The automatic method can be used with either closed-source or open-source MLLMs without requiring groundtruth segmentation. It can be used as an interpretability tool to have a better understanding of where MLLMs look and when the correct grounding occurs with respect to the output tokens. In the oracle, we rely on the ground-truth mask to select the correct token and its corresponding segmentation with the highest intersection over union as an upper bound. More details are provided in the supplementary.
Figure 4 shows qualitative examples from our interpretability tool, the oracle variant, applied to LLaVA 1.5, where the best masks and their respective noun phrases do not necessarily correspond to the queried referring expression. We also study the concept categories for the noun phrases where the best grounding occurs, relying on our interpretability tool, refer to Sec. 4.3. Figure 5 shows how the output from our interpretability mechanism and the oracle variant is better capable of identifying the fine-grained detail, “dorsal fin”, from LLaVA 1.5 without being trained with mask supervision, unlike other pixel-level MLLMs, while still maintaining its language capabilities. Considering that pixel-level MLLMs rely on standard ones as their base MLLM, our benchmarking and interpretability study questions whether there could be better pixel-level MLLMs built without the loss of their original language and grounding capabilities.
| Method | Venue | Grounding Output | PixMMVP | PixMMVP Prompt Sensitivity | |||||||
| w/ GPT | |||||||||||
| Pixel-level MLLMs w/ Mask Output | |||||||||||
| OMG LLaVA (7B) | NeurIPS 2024 | Mask | 12.0 | 12.0 | 17.8 (13.8) | 38.0 (38.8) | 18.2 (18.3) | 5.9 | 11.8 | 29.7 | 9.8 |
| GLAMM (7B) | CVPR 2024 | Mask | 1.3 | 2.7 | 31.5 (30.6) | 47.4 (48.3) | 5.1 (5.1) | - | - | 41.9 | - |
| LISA (7B) | CVPR 2024 | Mask | 7.3 | - | 18.1 (14.1) | 42.9 (42.9) | 12.5 (12.5) | - | - | 35.0 | - |
| LLaVA-G (7B) | ECCV 2024 | Mask | 9.3 | - | 19.2 (15.2) | 28.8 (29.1) | 14.0 (14.1) | - | - | 26.8 | - |
| RGA (7B) | ICCV 2025 | Mask | 5.3 | 51.3 | - | 52.7 (54.6) | 52.0 (52.9) | 23.2 | 28.5 | 45.8 | 35.1 |
| Standard MLLMs Extended for Mask Output | |||||||||||
| LLaVA 1.5 (7B) + SAM | baseline (ours) | Box2Mask | 27.3 | 28.0 | - | 38.2 (40.1) | 32.3 (33.0) | 18.4 | 18.7 | 21.9 | 20.2 |
| LLaVA 1.5 (13B) + SAM | baseline (ours) | Box2Mask | 39.3 | 30 | - | 45.5 (47.7) | 42.2 (43.1) | 24.4 | 23.7 | 24.9 | 24.6 |
| Cambrian (8B) + SAM | baseline (ours) | Box2Mask | 52.0 | 52.0 | - | 39.9 (41.8) | 45.2 (46.3) | - | 44.2 | 28.4 | 34.6 |
| Qwen2.5-VL (7B) + SAM | baseline (ours) | Point2Mask | 54.7 | 38.0 | - | 37.6 (36.3) | 44.6 (43.6) | 29.6 | 33.8 | 24.8 | 28.6 |
| Qwen2.5-VL (7B) + SAM | baseline (ours) | Box2Mask | 54.7 | 38.0 | - | 33.2 (31.0) | 41.3 (39.6) | 29.6 | 33.8 | 26.0 | 29.4 |
| LLaVA 1.5 (7B) + (a+s) | baseline (Arxiv 2025) | Point2Mask | 27.3 | 28.0 | 11.8 (7.8) | 18.1 (14.1) | 22.0 (18.8) | 18.4 | 18.7 | 16.5 | 17.5 |
| LLaVA 1.5 (13B) + (a+s) | baseline (Arxiv 2025) | Point2Mask | 39.3 | 30 | 12.1 (8.5) | 17.9 (13.9) | 24.6 (20.5) | 24.4 | 23.7 | 16.3 | 19.5 |
| Cambrian (8B) + (a+s) | baseline (Arxiv 2025) | Point2Mask | 52.0 | 52.0 | 19.5 (15.5) | 16.7 (13.3) | 28.4 (23.9) | - | 44.2 | 18.8 | 26.4 |
| LLaVA 1.5 (7B) + PixFoundation | interp. tool (ours) | Point2Mask | 27.3 | 28.0 | 31.2 (29.2) | 42.8 (42.1) | 33.9 (33.6) | 18.4 | 18.7 | 37.9 | 25.0 |
| LLaVA 1.5 (13B) + PixFoundation | interp. tool (ours) | Point2Mask | 39.3 | 30 | 27.1 (26.0) | 41.3 (40.1) | 40.3 (39.7) | 24.4 | 23.7 | 37.5 | 29.6 |
| Cambrian (8B) + PixFoundation | interp. tool (ours) | Point2Mask | 52.0 | 52.0 | 47.0 (44.7) | 41.8 (41.4) | 49.4 (48.1) | - | 44.2 | 33.6 | 38.2 |
| Qwen2.5-VL (7B) + PixFoundation | interp. tool (ours) | Point2Mask | 54.7 | 38.0 | - | 46.5 (44.2) | 50.3 (48.9) | 29.6 | 33.8 | 31.6 | 32.7 |
| LLaVA 1.5 (7B) + PixFoundation | upper bound (ours) | Point2Mask | 27.3 | 28.0 | 37.8 (34.8) | 55.3 (53.1) | 37.2 (36.7) | 18.4 | 18.7 | 56.4 | 28.1 |
| LLaVA 1.5 (13B) + PixFoundation | upper bound (ours) | Point2Mask | 39.3 | 30 | 36.8 (33.7) | 55.2 (53.0) | 45.9 (45.1) | 24.4 | 23.7 | 55.2 | 33.8 |
| Cambrian (8B) + PixFoundation | upper bound (ours) | Point2Mask | 52.0 | 52.0 | 68.7 (67.1) | 72.1 (70.8) | 60.4 (60.0) | - | 44.2 | 70.9 | 54.5 |
| Qwen2.5-VL (7B) + PixFoundation | upper bound (ours) | Point2Mask | 54.7 | 38.0 | - | 55.0 (52.8) | 54.8 (53.7) | 29.6 | 33.8 | 38.9 | 36.2 |
| Method | Grounding Output | PixCV-Bench | |||
| Pixel-level MLLMs w/ Mask Output | |||||
| OMG LLaVA (7B) | Mask | 12.0 | 42.1 | 50.5 | 45.9 |
| GLAMM (7B) | Mask | - | - | 51.9 | - |
| LISA (7B) | Mask | 3.7 | - | 48.1 | 6.7 |
| LLaVA-G (7B) | Mask | 14.1 | 4.4 | 17.6 | 15.8 |
| RGA (7B) | Mask | - | 72.6 | 51.1 | 56.0 |
| Standard MLLMs Extended for Mask Output | |||||
| LLaVA 1.5 (7B) + SAM | Box2Mask | 17.4 | 60.3 | 29.8 | 39.9 |
| LLaVA 1.5 (13B) + SAM | Box2Mask | 14.5 | 61.4 | 32.9 | 42.8 |
| Cambrian (8B) + SAM | Box2Mask | 62.2 | 72.2 | 34.2 | 46.4 |
| Qwen2.5-VL (7B) + SAM | Point2Mask | 48.9 | 60.8 | 33.7 | 43.4 |
| Qwen2.5-VL (7B) + SAM | Box2Mask | 48.9 | 60.8 | 48.9 | 54.2 |
| LLaVA 1.5 (7B) + (a+s) | Point2Mask | 17.4 | 60.3 | 15.7 | 24.9 |
| LLaVA 1.5 (13B) + (a+s) | Point2Mask | 14.5 | 61.4 | 14.9 | 24.0 |
| Cambrian (8B) + (a+s) | Point2Mask | 62.2 | 72.2 | 15.9 | 26.1 |
| LLaVA 1.5 (7B) + PixFoundation | Point2Mask | 17.4 | 60.3 | 34.1 | 43.6 |
| LLaVA 1.5 (13B) + PixFoundation | Point2Mask | 14.5 | 61.4 | 33.9 | 43.7 |
| Cambrian (8B) + PixFoundation | Point2Mask | 62.2 | 72.2 | 38.4 | 50.1 |
| LLaVA 1.5 (7B) + PixFoundation | Point2Mask | 17.4 | 60.3 | 49.7 | 54.5 |
| LLaVA 1.5 (13B) + PixFoundation | Point2Mask | 14.5 | 61.4 | 50.6 | 55.5 |
| Cambrian (8B) + PixFoundation | Point2Mask | 62.2 | 72.2 | 64.4 | 68.1 |
| Qwen2.5-VL (7B) + PixFoundation | Point2Mask | 48.9 | 60.8 | 43.6 | 50.8 |
@@ 4.1 Experimental setup
Evaluation benchmarks, protocols and metrics. PixMMVP is composed of 300 images paired with questions, choices, referring expressions and segmentation masks, while PixCV-Bench has 1,438 images with their corresponding annotations similarly. On each benchmark, we evaluate the VQA and referring segmentation following three probing techniques and report their metrics. The first probing is to evaluate the VQA ability, where the accuracy is computed using GPT-4o following [24] as, . If the model generates a segmentation without explicitly asking it to, it is evaluated with respect to the ground-truth referring segmentation in terms of mean intersection over union as . The second probing prompts the model to identify the referred expression and evaluates the mean intersection over union reported as . The third probing following [23] instructs the model to generate an option letter and evaluate the accuracy directly without GPT-4o, reported as . We evaluate the score of each model, , which is the harmonic mean across both referring segmentation and VQA,
For the language prompt sensitivity evaluation, we randomly generate 10 variations for each of the three language variants, and for the visual one, we generate four for each of the two variants. As such, we have in total 30 variations for the VQA and eight for referring segmentation; for simplicity, we report the mean of the metric used across all the variations. For the VQA accuracy, we follow the third probing, where accuracy is evaluated using the generated option’s letter with simple postprocessing on the model’s output. To avoid observations from previous work [9] that prompt sensitivity failures can stem from the evaluation and post-processing itself, we evaluate with GPT-4o the same output and report both, as and w/ GPT, respectively. We provide in the supplementary materials a language prompt sensitivity evaluation for the grounding aside from the VQA.
Compared methods. We focus on evaluating five state-of-the-art pixel-level MLLMs; LISA [12], GLAMM [20], OMG-LLaVA [34], LLaVA-G [33] and RGA [26]. These were either equipped with a SAM or SAM 2 decoder or with other powerful decoders, e.g., OMG-Seg. Some of these were trained with the SAM dataset, SA-1B, extended to the visual grounding [20] or other large-scale segmentation data. Furthermore, we evaluate the attend and segment (a+s) [3], the oracle relying on the highest intersection over union in mask selection (PixFoundation), the automatic selection (PixFoundation) and a simple baseline that uses the output bounding box/point to prompt SAM without any selection. These are implemented on top of four base MLLMs, which are LLaVA 1.5 (7B, 13B) [14], Cambrian (8B) [23] and Qwen2.5-VL [1]. The automatic selection is implemented using GPT-5.1. Additional details are in the supplementary.
@@ 4.2 Are the current pixel-level MLLMs heading in the right direction?
In order to answer this, we evaluate each of these pixel-level MLLMs’ capabilities in VQA and referring segmentation in challenging tasks. Table 1 shows the results on the challenging PixMMVP and the respective prompt sensitivity ones. From the accuracy of VQA, standard MLLMs that are not trained for mask output surpass their pixel-level counterpart with 3-14%. This is especially evident in the language prompt sensitivity, where RGA, the best pixel-level MLLM, drops VQA accuracy to 28.5%, which is worse than other standard MLLMs, i.e., Cambrian (8B) and Qwen2.5-VL (7B) at 44.2% and 33.8%, respectively. Furthermore, RGA shows severe failure in , which is attributed to its sensitivity to the VQA prompt, as it can only answer the multiple-choice question when instructed to generate a letter, while exhibiting limited ability to generate free-form answers. As for the referring segmentation, some of these pixel-level MLLMs, e.g., OMG-LLaVA and LLaVA-G, underperform the baselines.
Comparing PixFoundation and the baseline (a+s) clearly shows that our interpretability tool is better because of the understanding that the grounding does not necessarily coincide with the output noun phrase closest to the referring expression, unlike the (a+s) baseline. This is further confirmed in the point accuracy results in Table 3. When looking at the overall score, , in PixMMVP, our concurrent work (RGA) shows strong performance that is better than its respective model Qwen2.5-VL, yet in the prompt sensitivity analysis, it shows lower performance than Cambrian with PixFoundation. The second-best pixel-level MLLM, OMG-LLaVA, which was prior to the release of our benchmark, consistently lags behind our baselines.
Various factors can impact the language and visual abilities of these MLLMs, especially in visual grounding and VQA. These results could be attributed to training recipes, design choices or datasets; this is evident with RGA performance surpassing other pixel-level MLLMs, which rely on mixing both QA and segmentation datasets during training and expanding them with video tasks. However, even with RGA, its VQA abilities downgraded drastically in the language prompt sensitivity with a factor of 1.8, while the standard MLLM Cambrian degraded with a factor of 1.2. Unlike RGA, Qwen2.5-VL and Cambrian generate grounding output directly in language. There could be potential opportunities to improve the language capabilities through exploring these design choices further in the segmentation domain. Table 2 shows PixCV-Bench results. Our concurrent work, RGA, shows strong performance, followed closely by Qwen2.5-VL. It also shows that Point2Mask is worse than Box2Mask in our simple baselines in PixCVBench, which could be due to the referring expressions inquiring about multiple objects/instances more than in PixMMVP due to the nature of the questions, e.g., relative positioning. Furthermore, Point2Mask shows higher sensitivity to visual prompt variations than Box2Mask.
Summary. In summary, the state-of-the-art pixel-level MLLMs prior to the release of our benchmark have shown lower performance than simple baselines, where they specifically lag in their language abilities and visual question answering. Prior to our work, pixel-level MLLMs were mostly ignoring the language abilities, yet we show various examples that emphasize the role of language in visual grounding. Moreover, we show that visual grounding might not coincide with the output noun phrase most similar to the referred expression, where our oracle upper bound and automatic method both surpass the attend and segment. Our interpretability tool enabled such an important insight to help understand these MLLMs better.
| Base MLLM | PixFoundation | a+s | PixFoundation |
| LLaVA 1.5 (7B) | 58.7 | 24.3 | 51.4 |
| LLaVA 1.5 (13B) | 66.0 | 28.3 | 58.7 |
@@ 4.3 When does grounding emerge in MLLMs?
When - location. Taking into account the powerful performance of the oracle upper bound, it begs the question of when grounding emerges. We start by looking at when it emerges in terms of the location. We analyze the word/phrase location with respect to the full output text in terms of a percentage of its total length (i.e., 0% means the beginning of the text). Accordingly, Fig. 6(a) shows the location percentages histogram, binned at 10%, for the three base MLLMs reporting the oracle selection and evaluating on PixMMVP benchmark using the second probing. In the LLaVA 1.5 variants, the highest grounding is at the last 40%, while for Cambrian it is at the last 60%.
When - concept. For the second analysis, we look into the concept category that the correct output word/phrase corresponds to. The previous assumption, in other works is that grounding emerges in the exact noun/noun phrase of the object of interest. Except our analysis confirms that this is not necessarily the case. We take the correct noun/noun phrase where the grounding emerges based on the oracle from all three variants, then we pass it to GPT-4o to request a grouping of these concepts. It results in six main groups, which are: (i) color and appearance, (ii) location and position, (iii) object parts, (iv) context and setting, (v) objects and entities, and (vi) State. We then prompt for each of the noun/noun phrases, GPT-4o, to categorize it within these six categories. The histogram of the occurrences of these concept categories is shown in Fig. 6(b). It conveys that in certain scenarios, the correct output when grounding emerges can be describing the position or the color of the object. Figure 4 shows qualitative examples of these scenarios.
Summary. In summary, we found that emergent pixel-level grounding might not coincide with the input referring expression. We show that grounding in MLLMs can emerge in the noun phrase that corresponds to color, position or other characteristics of the object of interest.
@@ 5 Conclusion
We propose two benchmarks showing that pixel-level MLLMs degrade the ability in VQA and even grounding of fine-grained objects. We propose language and visual prompt variations tailored for visual grounding. Additionally, we provide an interpretability mechanism that showed an interesting insight on when visual grounding occurs in MLLMs. Our paired benchmarks and evaluation pave the road towards better and interpretable pixel-level MLLMs.
@@ Appendix A Additional implementation details
In this section, we cover additional details about our proposed datasets, the implementation of the evaluation setup, prompt sensitivity evaluation and baselines.
@@
Datasets. Our proposed datasets, PixMMVP and PixCV-Bench, are composed of referring expressions describing the object of interest in the respective question and its segmentation mask. We show in Fig. 8 examples of these ground-truth annotations for both datasets. It shows the challenging scenarios in pixel-level visual grounding, which is strongly tied to the visual question answering task, since an integral part of answering these questions requires the grounding of the object/s of interest.
Models. We also detail the model checkpoints we use for the four pixel-level MLLMs and their variants, retrieved from HuggingFace [28] in Table 4. These also include the model checkpoints used for the base MLLMs that were not trained with mask supervision. Furthermore, we provide details on the oracle selection mechanism. We discard the cases where the ground-truth segmentation is all background in the oracle selection in the when analysis and the point accuracy evaluation, since there is no ground-truth grounding to evaluate against. While in the quantitative and qualitative evaluation, we use an empty mask for the oracle. For the automatic variant, we automatically identify the images that do not have the referring expression, and correspondingly output an empty mask. Specifically, we prompt the models with the question: “Does this image have EXP? Answer with Yes/No only.” Similarly, for the attend and segment architecture (a+s), we follow their approach in filtering out the selected noun phrases that have a similarity less than 0.7 in the SpaCy embeddings and instead output an empty mask. This pre-processing is only needed in PixMMVP, while on PixCV-Bench we do not need such filtration across the (a+s), automatic or oracle variants. In the oracle upper bound, whenever the model is to be evaluated in a multiple object scenario, we take all the possible pairs of the masks and select the best pair based on the highest intersection over union.